# Distillation Improves Visual Place Recognition for Low Quality Images

### v2

@@ Abstract
Real-time visual localization often utilizes online computing, for which query images or videos are transmitted to remote servers for visual place recognition (VPR). However, limited network bandwidth necessitates image-quality reduction and thus the degradation of global image descriptors, reducing VPR accuracy. We address this issue at the descriptor extraction level with a knowledge-distillation methodology that learns feature representations from high-quality images to extract more discriminative descriptors from low-quality images. Our approach includes the Inter-channel Correlation Knowledge Distillation (ICKD) loss, Mean Squared Error (MSE) loss, and Triplet loss. We validate the proposed losses on multiple VPR methods and datasets subjected to JPEG compression, resolution reduction, and video quantization. We obtain significant improvements in VPR recall rates under all three tested modalities of lowered image quality. Furthermore, we fill a gap in VPR literature on video-based data and its influence on VPR performance. This work contributes to more reliable place recognition in resource-constrained environments. Source code and data available at https://github.com/ai4ce/LoQI-VPR.
@@ I Introduction
Visual Place Recognition (VPR) has rapidly become a foundational component in machine perception. By condensing RGB images into compact vector representations, devices can recognize previously visited locations and recall their locations through only vision input and an image database of known locations [1, 2]. VPR has become a critical component in applications ranging from robots navigating complex industrial settings to augmented reality systems overlaying digital data onto the physical world to wayfinding applications that support vulnerable populations [3, 4, 5].
The adaptability of VPR to diverse environments has expanded its use cases. From smartphones to autonomous vehicles, the need for efficient and precise VPR systems is rising. However, this demand, especially in consumer devices, introduces novel challenges. A primary concern is the rapid growth in scene diversity and volume for real-world VPR applications, leading to large image databases that pose storage challenges for devices with limited capacity [6, 7].
Furthermore, the goal of more precise image descriptor vectors and thus better VPR performance has led to more complex VPR methods [8, 9, 10]. While such advancements do not delay descriptor matching, they increase the computational demands of extracting descriptors from images. Thus, real-time VPR on standard devices becomes more challenging due to significant computational cost [11] in addition to storage needs.
@@
Figure 1: Enhanced Low Quality Image Retrieval: After knowledge distillation from high quality images, multiple VPR methods succeeds in their VPR recall of this JPEG-compressed query image from the Mapillary SLS dataset, while the same methods fail to do so beforehand.
@@
Transitioning to online computation, where images are transmitted over the network for centralized processing, could be a solution for both issues but not without difficulties. While many devices simultaneously capture images to query the VPR image database, real-time communication with servers becomes a bottleneck. As most real-world VPR applications localizes over image sequences, the underlying video streaming service could alleviate network bandwidth pressure through reducing resolution and increasing compression of images and quantization of video streams. However, these measures degrade image descriptors and reduce localization accuracy [12].
For this online VPR paradigm, we propose a VPR-specific knowledge-distillation framework that improves descriptor generation and thus VPR performance under low-quality images, offering the following contributions:
  • We develop a strategy to distill state-of-the-art VPR methods’ knowledge of high-quality images into that of low-quality images, significantly improving performance on low-quality images.
  • We demonstrate three distillation loss functions on multiple public datasets and analyze the nature of each on descriptor quality.
  • We validate the generalizability of our approach by extending it to multiple forms of image quality reduction, including resolution decrease and video quantization, curating a video-based VPR dataset in the process to also fill a gap in video data for VPR research.
@@ II Related Work
II-A VPR’s Challenges with Low-Quality Images
Prevailing VPR literature frames the problem of image-based localization as an image retrieval task [2]. Under this modality, VPR performance is heavily dependent on image descriptors’ abilities to relate images with distinct appearances to geographical proximity. Earlier VPR techniques such as GIST [13], HOG [14], and VLAD [15] generate compact image descriptors through hand-crafted local feature extractors. For better image matching performance, NetVLAD [8] and other recent VPR methods such as Patch-NetVLAD [16], MixVPR [9], AnyLoc [10], DINOv2 SALAD [17], Bag-of-Queries [18], and CricaVPR [19] leverage progress in deep learning and utilize convolutional and attention-based neural networks to produce learned local features and extract global image-level descriptors.
However, the rising complexity of VPR global descriptor extraction, while improving image retrieval accuracy, poses computational challenges, particularly in resource-constrained devices across real-world applications. Although special-purpose VPR techniques [11, 20, 21] with low computational cost address this limitation, the transmission of images for centralized descriptor extraction is often still necessary for deploying VPR techniques at scale [22, 23], completely free of both storage and computational constraints. Nevertheless, the resulting decrease in image quality from networking limitations reduces localization accuracy, which occurs under both lowered image resolution [12] and increased JPEG compression [24]. As many existing VPR datasets consist of purely JPEG images [25, 26, 27, 28, 29, 30, 31, 32, 33], the JPEG format’s prevalence both in VPR and in general [34] makes JPEG compression a main avenue of image quality reduction, besides decreasing image resolution. To our best knowledge, there exists a gap in VPR literature on publicly available datasets offering video data, and the effects of video quantization on VPR performance remain largely unexplored.
@@ II-B Knowledge Distillation and Transfer Learning in VPR
Likewise, research on effective methods to extract more representative global descriptors from low-quality images is also lacking in the VPR community. Existing solutions have addressed the issue at the descriptor matching and image-retrieval level. The authors of [24] propose a sequence-based image-retrieval strategy [35], which queries the database images multiple times with lower per-query cost thanks to JPEG compression. Furthermore, predicting the retrieval success for each query image could favor points within the sequential queries that are more easily localized [36]. We instead wish to explore knowledge distillation as a means to directly improve descriptor extraction and thus increase the accuracy of database retrieval via individual queries.
@@ III Distillation Methodology
Our goal was to enhance existing VPR methods to extract more representative global descriptors when given lower-quality images. Thus, we apply knowledge distillation techniques at the descriptor extraction level of VPR. For any VPR method, we aim to allow a student descriptor extractor handling low-quality images to approximate the output of its teacher counterpart, handling high-quality images.
Conceptual Model of Distillation for VPR
Given some VPR method, we initialize it twice to form the teacher branch $f_{t}$ and student branch $f_{s}$, processing high-quality image $I^{h}$ and its low-quality counterpart $I^{l}$ respectively. As illustrated by fig. 2, we conceptualize the teacher model as consisting of a feature encoder $f^{\theta}_{t}$ and descriptor aggregator $g^{\phi}_{t}$. Likewise, the student consists of a structurally identical encoder $f^{\theta}_{s}$ and aggregator $g^{\phi}_{s}$. Between the encoder and aggregator, the student and teacher branch generates 3D latent code $z^{l}$ and $z^{h}$, respectively. Finally, both branches aggregate their latent codes into global descriptors $v^{l}$ and $v^{h}$. Our distillation methodology has three loss functions that operate on this understanding of VPR methods.
ICKD Loss
The Inter-Channel Correlation Knowledge Distillation (ICKD) loss [40] operates on the latent codes $z^{l}$ and $z^{h}$. Given the potential size differences between inputs $I^{l}$ and $I^{h}$, the latent codes $z^{l}$ and $z^{h}$ may differ in dimensions. Traditional distillation losses, which typically require matching dimensions for the latent codes, are not directly applicable. Instead, the ICKD loss computes a standard-size Inter-Channel Correlation (ICC) matrix for both $f^{\theta}_{s}$ and $f^{\theta}_{t}$, allowing the former to approximate the latter.
To illustrate ICKD loss, define $I^{l}$ with resolution $W\times H$ and let the dimensions of their latent codes $z^{h}$ and $z^{l}$ be $c\times W^{\prime}\times H^{\prime}$ and $c\times w^{\prime}\times h^{\prime}$, respectively. Leveraging $c$ being fixed, the ICKD concatenates along the remaining dimensions $w^{\prime}$ and $h^{\prime}$ to produce vector $p^{l}$ with dimensions $c\times w^{\prime}\cdot h^{\prime}$. $p^{l}$ is then normalized along its second dimension, producing $\hat{p^{l}}=\frac{p^{l}}{\|p^{l}\|}$. The subsequent ICC matrix $C^{l}=\hat{p^{l}}\times\hat{p^{l}}^{T}$ with dimensions $c\times c$ is normalized to $\hat{C^{l}}=\frac{C^{l}}{\|C^{l}\|}$. A similar procedure for $z^{h}$ yields $\hat{C^{h}}\in\mathbb{R}^{c\times c}$. The ICKD loss, formulated to ensure $f^{\theta}_{s}$ learns from $f^{\theta}_{t}$, is defined as:
@@
Mean Squared Error (MSE) Loss
In addition to comparing latent codes, we directly compute the discrepancies between the image descriptors $v^{l}$ and $v^{h}$. The MSE loss is expressed as:
@@
The MSE loss trains $f_{s}$ to produce similar global descriptors as $f_{t}$ despite different input images, gaining insights from $f_{t}$ by directly emulating $f_{t}$’s outputs [42].
Weakly Supervised Triplet Ranking Loss
Applied exclusively within the student branch $f_{s}$, this loss is inspired by NetVLAD [8]. It ensures that the input image’s descriptor $v^{l}$ is more similar to that of a geographically proximate (positive) image, $\{v^{l}_{p}\}$, than that of a distant (negative) image, $\{v^{l}_{n}\}$. The goal is to satisfy the inequality $d_{\theta}(v^{l},v^{l}_{p})<d_{\theta}(v^{l},v^{l}_{n})$, where $d_{\theta}$ denotes the Euclidean distance between descriptors.
The triplet loss is articulated as:
$L_{\theta 3}=\sum_{j}l\left(\min_{i}d_{\theta}^{2}(v^{l},{v^{l}_{p}}_{i})-d_{\theta}^{2}(v^{l},{v^{l}_{n}}_{j})+m\right).$ (3)
In this equation, $l$ represents the hinge loss, defined as $l(x)=\max(x,0)$. The term $m$ is a margin parameter, ensuring a buffer between positive and negative pairs, thereby enhancing the discriminative capabilities of the learned embeddings. The component $\min_{i}d_{\theta}^{2}(v^{l},{v^{l}_{p}}_{i})$ selects the descriptor of the most similar image as estimated by the student model $f_{s}$, ensuring the loss is attuned to the nearest matches.
Composite Loss
We also aim to assess the effectiveness of a combined loss function relative to the individual application of the aforementioned losses. We compute a cumulative loss as:
$L=L_{\theta_{1}}+\alpha L_{\theta_{2}}+\beta L_{\theta_{3}}.$ (4)
In this equation, $\alpha$ and $\beta$ are weighting coefficients for $L_{\theta 2}$ and $L_{\theta 3}$, balancing the influence of each loss within the composite function. Our experiments will also examine each loss’s individual effects on distillation.
IV Experimental Setup
To analyze the efficacy of $L_{\theta 1}$, $L_{\theta 2}$, $L_{\theta 3}$, and $L$, we performed distillation on the student branch $f_{s}$ using the GSV-Cities VPR dataset [32]. Afterwards, we evaluated $f_{s}$’s retrieval accuracy via the Deep Visual Geo-localization Benchmark [43] and some of its available testing datasets. For the following training and testing datasets, we define low quality images $I^{l}$ as their original images ($I^{h}$) after $90$% JPEG compression (quality level = $10$).
Following Tomita et al. [35], we evaluate $f_{s}$ with $I^{l}$ VPR database and queries to avoid arbitrarily hampering descriptor matching from inconsistent compression levels, which also confirms the advantage of decreased database building cost when expecting $I^{l}$ queries. Here, we reduce image quality via JPEG compression due to the absence of video data and other image formats. We would explore our methodology’s generalizability to low image resolution and high video quantization in section VI.
@@ IV-A1 Training Dataset
The GSV-Cities dataset contains approximately $530,000$ street view images of cities around the world collected from various years. The $23$ total cities in the dataset are divided into a higher and lower resolution group containing images at $640\times 480$ and $400\times 300$. To decrease training cost, we designated only the cities with lower resolution images to be used, effectively shrinking the dataset to approximately $194,000$ images. This selective use of GSV-Cities additionally investigates whether knowledge distillation could be successful without high quality training data in the absolute sense.
Table I: Image Count of Testing Datasets
Partition Mapillary SLS Nordland Tokyo 24/7
Database 18871 27592 75984
Queries 11084 2760 315

### v3

@@ Abstract
Real-world applications of Visual Place Recognition (VPR) often rely on cloud computing, where query images or videos are transmitted elsewhere for visual localization. However, limited network bandwidth forces a reduction in image quality, which degrades global image descriptors and consequently VPR accuracy. We address this issue at the descriptor extraction level with a knowledge-distillation framework that transfers feature representations from high-quality images to low-quality ones, allowing VPR methods to produce more discriminative descriptors from the latter. Our approach leverages three complementary loss functions —- the Inter-channel Correlation Knowledge Distillation (ICKD) loss, Mean Squared Error (MSE) loss, and a weakly supervised Triplet loss —– to guide a student network in approximating the high-quality features produced by a teacher network. Extensive experiments on multiple VPR methods and datasets subjected to JPEG compression, resolution reduction, and video quantization demonstrate significant improvements in recall rates. Furthermore, our work fills a gap in the literature by exploring the impact of video-based data on VPR performance, thereby contributing to more reliable place recognition in resource-constrained environments. Source code and data are available at https://ai4ce.github.io/LoQI-VPR/.
@@ I Introduction
Visual Place Recognition (VPR) has rapidly become a foundational component in machine perception. Through learning compact vector representations of RGB images, or global image descriptors, VPR methods localize unseen images by recalling one or more known locations among a database of images based on descriptor similarity [1, 2]. VPR is critical for real-world applications ranging from robotic navigation in complex industrial environments to augmented reality systems and way finding aids for vulnerable populations [3, 4, 5].
However, VPR deployment at scale comes with an inherent dilemma between VPR performance and computational / networking costs. On one hand, state-of-the-art (SOTA) VPR methods leverage deep learning to produce highly discriminative and generalizable global descriptors [6, 7]. However, such methods are computationally costly for both descriptor extraction and matching [8], making large-scale deployment on consumer / edge devices impractical. Besides computational cost, there is also a further dilemma caused by the storage demands of the VPR image database, which is difficult to maintain locally given the rapid growth in scene diversity and volume for real-world VPR applications [9, 10].
@@
Figure 1: Enhanced Low Quality Image Retrieval: After knowledge distillation from high quality images, this JPEG-compressed query image from the Mapillary SLS dataset (top center) is correctly localized by multiple VPR methods (top row), whereas they failed beforehand (bottom row).
@@
Given these challenges, one option is to perform VPR locally on low-end devices using methods specifically designed for low computational cost. The second option is to perform VPR over the cloud, where images are transmitted elsewhere for centralized processing. Arguably, the cloud VPR approach has greater potential for accurate localizations, as it carries the aforementioned benefits of SOTA VPR methods. However, network bandwidth constraints are common when many devices simultaneously capture images to query the VPR database, hampering real-time communication with servers. As most real-world VPR applications localize over image sequences, the underlying video streaming service could alleviate bandwidth pressures by reducing resolution and increasing image or video compression. However, such quality reductions to VPR methods’ visual inputs significantly hamper localization accuracy [11].
Our work offers a potential solution to the dilemmas of real-world VPR, benefiting from the performance and generalizability of SOTA VPR methods while mitigating VPR accuracy loss from lowered image quality. As networking-induced image quality reduction within the cloud VPR paradigm is not easily eliminated, we propose a knowledge distillation framework to improve the performance of VPR methods under low-quality images. Although image processing techniques not specific to VPR such as super-resolution [12, 13] and image deblurring [14, 15] can potentially restore images’ visual contents after network transmission, these methods do not account for images’ role as place representations and thus are not directly beneficial for VPR performance. In short, we offer the following contributions:
  • Utilizing knowledge distillation, we develop a training framework to improve the localization performance of VPR global descriptors extracted from low-quality images. Our method is not limited to any particular VPR method.
  • We validate our framework across multiple public VPR datasets and various types of low-quality images, including JPEG compression, resolution reduction, and video quantization. In doing so, we also curate a video-based VPR dataset to address the low availability of video data for VPR research.
@@ II Related Work
II-A VPR Under Low-Quality Image Conditions
Prevailing VPR literature frames the problem of image-based localization as an image retrieval task, and research in VPR methods has traditionally focused on improving retrieval accuracy under high-level visual changes such as illumination changes, seasonal variation, and viewpoint shifts [2]. However, the impact of image quality degradation (e.g., low resolution, heavy compression, or motion blur) has received comparatively less attention. Recent studies have begun to highlight this gap. For instance, the accuracy of deep learning-based VPR methods such as NetVLAD [16] suffers significantly from low-resolution images [11]. In addition to lowered image resolution, increased JPEG compression has also been shown to negatively impact VPR methods [17]. Given the high compatibility and widespread adoption of the JPEG image format [18], it is prevalent in image transmission when deploying VPR at scale [19]. Furthermore, many widely used VPR datasets for studying the aforementioned high-level retrieval challenges also store images in JPEG format. Examples include Tokyo247 [20] (illumination changes), Mapillary SLS [21] (large temporal variation), Nordland [22] (seasonal variation), and GSV-Cities [23] (viewpoint shifts). Although in our work we consider multiple forms of image quality reduction, we place a special emphasis on JPEG compression for its relevance to real-world deployments and existing VPR datasets.
Besides image quality reduction, the cloud VPR paradigm also faces the challenge of video compression. Within VPR research, localization accuracy under video compression and specifically video quantization is not extensively studied. Even lightweight VPR methods [8, 24, 25], the alternative solution to the dilemmas of computational cost / storage demands introduced in I, still focus on the image domain. As video quantization introduces additional artifacts that further degrade image quality, we seek to better understand video compression’s effects on VPR descriptor extraction. To our best knowledge, we also contribute towards a literature gap on publicly available datasets offering video data.
@@ II-B Knowledge Distillation and Transfer Learning in VPR
Likewise, research on effective methods to extract more representative global descriptors from low-quality images is lacking in the VPR community. Existing solutions have addressed the issue at the descriptor matching and image-retrieval level. The authors of [17] propose a sequence-based image-retrieval strategy [26], which queries the database images multiple times with lower per-query cost thanks to JPEG compression. Furthermore, predicting the retrieval success for each query image could favor points within the sequential queries that are more easily localized [27]. We instead wish to explore knowledge distillation as a means to directly improve descriptor extraction and thus increase the accuracy of database retrieval via individual queries.
@@ III Distillation Methodology
Our goal is to enhance existing VPR methods by enabling them to extract more representative global descriptors from low-quality images. To achieve this, we apply knowledge distillation techniques directly at the descriptor extraction level. Specifically, we enable a student descriptor extractor, which processes low-quality images, to approximate the outputs of a teacher network operating on high-quality images. This approach leverages the rich, discriminative information from high-quality images to compensate for the information loss in degraded inputs.
III-A Conceptual Model of Distillation for VPR
Given some VPR method, we initialize two copies, designating one as the teacher branch $f_{t}$ and the other as the student branch $f_{s}$. The teacher processes a high-quality image $I^{h}$ while the student processes its low-quality counterpart $I^{l}$. As illustrated in Fig. 2, the teacher model comprises a feature encoder $f^{\theta}_{t}$ and a descriptor aggregator $g^{\phi}_{t}$. Similarly, the student consists of a structurally identical encoder $f^{\theta}_{s}$ and aggregator $g^{\phi}_{s}$. Between the encoder and aggregator, the teacher and student produce intermediate 3D latent codes $z^{h}$ and $z^{l}$, respectively. These latent codes are subsequently aggregated into global descriptors $v^{h}$ and $v^{l}$. Our distillation strategy is built on three distinct loss functions designed to transfer the high-quality feature representation from the teacher to the student.
III-B ICKD Loss
The Inter-Channel Correlation Knowledge Distillation (ICKD) loss [31] is employed to ensure that the student network’s latent code $z^{l}$ closely mimics the teacher network’s latent code $z^{h}$ in a channel-wise manner. In Visual Place Recognition (VPR), each channel of the latent code can be viewed as a local feature descriptor, and it is critical that these features remain consistent regardless of image quality. By enforcing similarity in the inter-channel correlations between $z^{l}$ and $z^{h}$, we help preserve the discriminative structural information necessary for robust localization.
Owing to potential differences in spatial dimensions between $I^{l}$ and $I^{h}$, the latent codes $z^{l}$ and $z^{h}$ may also differ in size, making traditional distillation losses inapplicable. To overcome this, the ICKD loss computes an Inter-Channel Correlation (ICC) matrix from the latent codes, effectively reducing their spatial dimensions to a consistent $c\times c$ form.
More specifically, let $I^{l}$ have resolution $W\times H$ and let the latent codes $z^{h}$ and $z^{l}$ have dimensions $c\times W^{\prime}\times H^{\prime}$ and $c\times w^{\prime}\times h^{\prime}$, respectively. By concatenating the spatial dimensions $w^{\prime}$ and $h^{\prime}$ for $z^{l}$, we form a vector $p^{l}$ of size $c\times(w^{\prime}\cdot h^{\prime})$. This vector is normalized along its second dimension to yield $\hat{p^{l}}=\frac{p^{l}}{\|p^{l}\|}$. The ICC matrix is then computed as $C^{l}=\hat{p^{l}}\hat{p^{l}}^{T}$ (with dimensions $c\times c$) and normalized to obtain $\hat{C^{l}}=\frac{C^{l}}{\|C^{l}\|}$. A similar process applied to $z^{h}$ results in $\hat{C^{h}}$. The ICKD loss is formulated as:
@@ III-B ICKD Loss
ensuring that the student encoder learns to mimic the channel correlations of the teacher, thereby preserving discriminative structural information despite spatial resolution differences.
III-C Mean Squared Error (MSE) Loss
To directly align the global descriptors produced by the teacher and student branches, we employ a Mean Squared Error (MSE) loss. This loss minimizes the difference between $v^{l}$ and $v^{h}$, enforcing that the student produces descriptors similar to those of the teacher despite the quality degradation of its input:
@@ III-C Mean Squared Error (MSE) Loss
This term provides a straightforward supervision signal, encouraging the student to closely emulate the teacher’s output.
III-D Weakly Supervised Triplet Ranking Loss
The weakly supervised Triplet Ranking Loss is applied solely within the student branch $f_{s}$ to further enhance the discriminative power of the learned descriptors. In our framework, following the same strategy as NetVLAD [16], positive and negative samples are defined based on geographic proximity: images captured within a predefined distance (e.g., 25 meters) of the query are treated as positives, while those farther away are considered negatives. Inspired by this framework, our loss ensures that the descriptor $v^{l}$ of a query image is closer (in Euclidean distance) to that of a geographically proximate (positive) image $v^{l}_{p}$ than to that of a distant (negative) image $v^{l}_{n}$. Formally, the loss enforces:
$d_{\theta}(v^{l},v^{l}_{p})<d_{\theta}(v^{l},v^{l}_{n}),$
where $d_{\theta}$ denotes the Euclidean distance.
The triplet loss is expressed as:
$L_{\theta 3}=\sum_{j}l\left(\min_{i}d_{\theta}^{2}(v^{l},{v^{l}_{p}}_{i})-d_{\theta}^{2}(v^{l},{v^{l}_{n}}_{j})+m\right),$ (3)
with $l(x)=\max(x,0)$ being the hinge loss and $m$ a margin parameter that ensures a sufficient gap between positive and negative pairs. Here, the term $\min_{i}d_{\theta}^{2}(v^{l},{v^{l}_{p}}_{i})$ selects, among a set of candidate positives in the batch, the descriptor that is closest to the query. This selection is performed based on the current descriptor distances, which serves as a weak supervision signal because it relies on geographic proximity and descriptor similarity rather than explicit ground-truth labels. Although this strategy may occasionally select a false positive, the margin $m$ and continuous network updates help mitigate this risk by ensuring that the model learns to focus on the most challenging and truly similar examples.
III-E Composite Loss
Finally, the overall distillation objective is defined as a composite loss combining the three individual terms:
$L=L_{\theta 1}+\alpha L_{\theta 2}+\beta L_{\theta 3},$ (4)
where $\alpha$ and $\beta$ are weighting coefficients that balance the influence of the MSE and Triplet losses relative to the ICKD loss. In our experiments, we analyze both the individual and combined effects of these losses to validate the efficacy of our distillation approach.
IV Experiments
To evaluate the effectiveness of our distillation methodology, we analyze the impact of each loss component ($L_{\theta 1}$, $L_{\theta 2}$, $L_{\theta 3}$) as well as their composite form $L$ on VPR performance. We perform distillation on the student branch $f_{s}$ using the GSV-Cities VPR dataset [23] and subsequently evaluate retrieval accuracy on several benchmark datasets from the Deep Visual Geo-localization Benchmark [33]. For all experiments, low-quality images $I^{l}$ are generated by applying 90% JPEG compression (quality level = 10) to the original high-quality images $I^{h}$. This choice allows us to simulate bandwidth-constrained scenarios while keeping the degradation consistent across training and testing.
Following the protocol in [26], we evaluate $f_{s}$ using a database and query set that both consist of $I^{l}$ images. This controlled setting avoids inconsistencies due to variable compression levels and also illustrates the reduced cost of database construction when operating under strict quality constraints. Although our current experiments focus on JPEG compression, we further explore the generalizability of our method to low image resolution and high video quantization in Section VI.
@@ IV-A1 Training Dataset
We use the GSV-Cities dataset, which contains approximately 530,000 street view images from 23 cities. To reduce training cost and assess our method under constrained data conditions, we use only the low-resolution subset ($400\times 300$), reducing the training set to about 194,000 images.